OpenStudy (anonymous):

what is the simplified form of (4x+2)/(x+5) + (3x+1)/(x+5)?

OpenStudy (anonymous):

7x + 3 4x + 3x - 2 + 5

OpenStudy (anonymous):

how on earth did you get that ?

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

4x-6x+10+(8x/3)-(14/3)=0 -2x+8x/3=2x/3 & 10+(-14/3)=16/3 when you add both 2x/3 + 16/3=0 2x+16=0 (multiplying the complete equation with 3) 2x=-16 x=-8 the answer can't be 8 its -8

OpenStudy (anonymous):

is that the answer ...?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

A) x+1 / x+5 , B) x+3 / x+5 , C) 7x+1 / x+5 , D) 7x+3 / x+5

OpenStudy (anonymous):

is that right ?

OpenStudy (anonymous):

And thanks for the medal :)

OpenStudy (anonymous):

idk i sent you the choices :P & no problem :)

OpenStudy (btaylor):

@alylee49 @justindrake has no clue what he's doing the denominators are the same, so just add the numerators. you get: \[\large{\frac{7x+3}{x+5}}\]

OpenStudy (anonymous):

@btaylor hah okay thanks :)

OpenStudy (anonymous):

Solve each equation. Check for extraneous solutions. Equation: |5x-1|+7=3x Here's what I did: |5x-1|+7=3x -7 -7 |5x+1|=3x-7 5x-1=3x-7 OR 5x-1=-(3x-7) +1 +1 5x=3x-6 -3x -3x 2x=-6 -- -- 2 2 x=-3 OR 5x-1=-(3x-7) 5x-1=-3x+7 8x = 8 x = 1 So, x=-3 or x=1. CHECK: |5x-1|+7=3x |5(-3)-1|+7=+3(-3) |-15-1|+7=-9 |-16|+7=-9 16+7=-9 23=9 But 23 doesn't equal 9. At first I thought it was just an exraneous solution. But then I checked the second number: |5x-1|+7=3x |5(1)-1|+7=3(1) |5-1|+7=3 |4|+7=3 4+7=3 11=3 But again, 11 doesn't equal 3. ----------------------------------------- Your work is correct. Both "solution" are extraneous. --- If you graph the left side and the right side of the equation separately, you will see that they do not intersect, i.e. they are never equal.

OpenStudy (anonymous):

here you go alye

OpenStudy (anonymous):

@BTaylor i have clue

OpenStudy (anonymous):

@justindrake i'm not looking for what x is though :P just the simplified form.

OpenStudy (anonymous):

@alylee49 but i helped u to

OpenStudy (btaylor):

@justindrake how can you solve if there's no equals sign?

OpenStudy (anonymous):

shut up bt please i dont want to fight with YOU BOO

OpenStudy (btaylor):

alright. I'll shut up. but i'm not even talking...